A solution at 25 degrees C is made by adding 0.095 g of calcium hydroxide (a soluble solid) into enough water to make a 175 mL solution.  What are the pH and pOH of the resulting solution?




1
Expert's answer
2021-01-15T06:31:04-0500

Ca(OH)2 = Ca2+ + 2OH-

So 1 mol calcium hydroxide give 2 mol OH-

Molar mass of Ca(OH)2 = 40 + (16×2) + (1×2) = 40 +32 +2 = 74 g mol-1

Mol of calcium hydroxide = (0.095\74)

= 0.00128 mol

So mol of OH- = 2 × 0.00128 = 0.00256 mol

So conc of OH- = (0.00256×1000) \ 175 = 0.0147

Using definition of molarity the above expression formed.

POH = -log[OH-] = -log 0.0147 = 1.83 (ans)

POH + PH = 14

So PH = 14 - POH = 14 - 1.83 = 12.17 ( ans)